### v1

# Inversion sequences growing on the left

Abstract
This work concerns a construction of pattern-avoiding inversion sequences from right to left, called a generating tree growing on the left. We first apply this construction to inversion sequences avoiding 201 and 210, resulting in a new way of computing their generating function. We then use a slightly modified construction to compute the generating function of inversion sequences avoiding 010 and 102, which was only conjectured before.
[…]
biblio.bib
Generating trees growing on the left for pattern-avoiding inversion sequences
Benjamin Testart [^1] Université de Lorraine, CNRS, Inria, LORIA, F 54000 Nancy, France
Keywords: pattern avoidance, inversion sequences, enumeration, generating trees, generating functions
1 Introduction
1.1 Basic definitions
For all $n\geqslant 0$, let $\mathcal{I}_{n}$ be the set of inversion sequences [Mansour_Shattuck_2015, Corteel_Martinez_Savage_Weselcouch_2016] of size $n$, that is the set of sequences $\sigma=(\sigma_{1},\dots,\sigma_{n})$ such that $\sigma_{i}\in\{0,\dots,i-1\}$ for all $i\in\{1,\dots,n\}$. Let $\mathcal{I}=\coprod_{n\geqslant 0}\mathcal{I}_{n}$ be the set of all inversion sequences.
We define a pattern to be a finite, nonempty sequence of nonnegative integers $\rho$ whose set of values forms an interval of integers starting at 0. In this work, we exclusively study patterns of length 3, which we denote $\rho_{1}\rho_{2}\rho_{3}$ instead of $(\rho_{1},\rho_{2},\rho_{3})$. For instance, $102$ and $010$ are patterns, but $022$ and $123$ are not.
We say that an inversion sequence $\sigma\in\mathcal{I}_{n}$ contains a pattern $\rho$ of length $k\leqslant n$ if there exists a subsequence of $k$ entries of $\sigma$ whose values appear in the same order as the values of $\rho$. Such a subsequence is called an occurrence of $\rho$. An inversion sequence avoids the pattern $\rho$ if it does not contain $\rho$. For instance, the sequence $(0,1,0,3,0,2,3)$ avoids the pattern 110, but it contains the pattern 021 since the subsequence $(0,3,2)$ is an occurrence of 021. If $P$ is a set of patterns, we denote by $\mathcal{I}(P)$ the set of inversion sequences avoiding all patterns in $P$, and we denote by $\mathcal{I}_{n}(P)$ its subset of sequences of size $n$.
1.2 Context and summary of results
Pattern-avoiding inversion sequences were introduced in [Mansour_Shattuck_2015] and [Corteel_Martinez_Savage_Weselcouch_2016], by analogy with pattern-avoiding permutations. Afterwards, many more articles studied inversion sequences avoiding patterns of length three, see [Martinez_Savage_2018, Beaton_Bouvel_Guerrini_Rinaldi_2019, Yan_Lin_2020, Kotsireas_Mansour_Yildirim_2024, Chen_Lin, Pantone201+210, Testart2024], among others.
Our work is concerned with two distinct classes of pattern-avoiding inversion sequences: $\mathcal{I}(201,210)$ and $\mathcal{I}(010,102)$. The class $\mathcal{I}(201,210)$ is counted by sequence A212198 in [OEIS]. It has an algebraic generating function, which was found in both [Chen_Lin] and [Pantone201+210], using two different constructions. Chen and Lin [Chen_Lin] use a decomposition around the rightmost saturated [^2] In an inversion sequence $\sigma$, the entry $\sigma_{i}$ is said to be ”saturated” if $\sigma_{i}=i-1$ (since $i-1$ is the largest value which could appear at position $i$). entry of inversion sequences, with one parameter counting the number of saturated entries. The construction used by Pantone [Pantone201+210] is a generating tree in one integer label and two boolean labels, and it involves some ”commitments” to future decisions.
The class $\mathcal{I}(010,102)$ is counted by sequence A374553 in [OEIS]. This sequence was computed in polynomial time in [Testart2024] using a decomposition involving $\{010,102\}$-avoiding words, and it was conjectured to have an algebraic generating function in [Pantone201+210] using the software [guessfunc].
In Section 2, we describe a construction for inversion sequences, which we call the ”generating tree growing on the left”. This construction first appeared in [Corteel_Martinez_Savage_Weselcouch_2016] to enumerate the classes $\mathcal{I}(101)$ and $\mathcal{I}(110)$. A modified construction based on this one also appears in [Lin_Fu_2021] to enumerate inversion sequences avoiding the vincular pattern $\underline{12}0$. By considering the generating tree growing on the left for inversion sequences avoiding 201 and 210, we find a new and simpler way of computing the generating function of $\mathcal{I}(201,210)$, in Section 3. We then modify this construction to apply it to inversion sequences avoiding 010 and 102 in Section 4. This allows us to compute the generating function of $\mathcal{I}(010,102)$ which was conjectured in [Pantone201+210].
As it is often the case with generating trees, we first turn the succession rule describing the tree into a functional equation for a multivariate generating function; we then solve this equation using the kernel method [GFforGT]. The equation we obtain for $\mathcal{I}(010,102)$ contains two evaluations $A(t,1)$ and $A(t,2)$ of a bivariate series $A$. This occurs because the succession rule produces labels with multiplicities that are powers of 2. In the case of $\mathcal{I}(201,210)$, the functional equation involves two catalytic variables, but it is easily solved by simply applying the kernel method twice.
1.3 Notation
For any integers $a,b\in\mathbb{Z}$, we denote the integer interval $[a,b]=\{k\in\mathbb{Z}\;:\;a\leqslant k\leqslant b\}$. We denote by $\varepsilon$ the empty sequence, that is the only sequence of size 00. We define $\chi$ to be the characteristic function of a statement, that is $\chi(\text{S})=\begin{cases}0&\text{if \; S is false},\\ 1&\text{if \; S is true}.\end{cases}$
For all $n\geqslant 0$ and $\sigma\in\mathcal{I}_{n}$, let
  • $|\sigma|=n$ be the size of $\sigma$,
  • $\mathbf{Vals}(\sigma)=\{\sigma_{i}\;:\;i\in[1,n]\}$ be the set of values of $\sigma$,
  • $\mathbf{min}(\sigma)=\mathbf{min}(\mathbf{Vals}(\sigma))$ be the minimum of $\sigma$, with the convention $\mathbf{min}(\varepsilon)=+\infty$,
  • $\mathbf{max}(\sigma)=\mathbf{max}(\mathbf{Vals}(\sigma))$ be the maximum of $\sigma$, with the convention $\mathbf{max}(\varepsilon)=-1$,
  • $\mathbf{Zeros}(\sigma)=\{i\in[1,n]\,:\,\sigma_{i}=0\}$ be the set of (positions of) zeros of $\sigma$.
2 Generating trees for inversion sequences growing on the left
A generating tree [West1, West2] is a rooted, labelled tree such that the label of each node determines its number of children, and their labels. A generating tree can be defined by a succession rule
$\Omega=\left\{\begin{array}[]{rclll}a\\ \ell&\leadsto&i^{m(\ell,i)}&\text{for}&i\in C(\ell)\\ \end{array}\right.$
composed of an axiom $a$, which labels the root of the tree, and a production which associates to each label $\ell$ the labels of the children of any node labelled $\ell$ in the tree. In the rule $\Omega$ above, each node labelled $\ell$ has $m(\ell,i)$ children labelled $i$, for each $i$ in a finite set of labels $C(\ell)$.
Generating trees are used to solve enumeration problems by identifying combinatorial objects with the nodes of a tree, so that each object of size $n$ corresponds to a node at level $n$ in the tree. This is useful when many nodes are roots of isomorphic subtrees, as the combinatorial objects corresponding to these nodes may all be given the same label. A succession rule for the generating tree then induces a recurrence relation on the number of objects of each size (which usually involves some parameters corresponding to the labels of the tree), or equivalently a functional equation for the (multivariate) generating function. For a more thorough introduction to generating trees in the context of pattern-avoiding inversion sequences, see [Testart2024, Section 2.1].
We define the generating tree for inversion sequences growing on the left as the only generating tree which has one node labelled by each inversion sequence (we now identify the nodes with their labels), whose root is the empty sequence $\varepsilon$, and in which the parent of each nonempty sequence is obtained by deleting its leftmost entry (which is always a zero), and subtracting 1 from all its positive entries. In this tree, the children of an inversion sequence are obtained by
  1. 1.
    first adding 1 to every positive entry in the sequence,
  2. 2.
    then adding 1 to some (or none) of the zero entries,
  3. 3.
    lastly, inserting a zero at the leftmost position.
Observe that this construction is indeed a generating tree (it produces only inversion sequences and, starting from the empty sequence $\varepsilon$, it produces every inversion sequence exactly once). Note that steps 1 and 3 are deterministic, whereas step 2 involves choosing a subset of the zero entries. For any inversion sequence $\sigma\in\mathcal{I}$, and any $Z\subseteq\mathbf{Zeros}(\sigma)$, we denote by $\mathbf{child}(\sigma,Z)=0\cdot(\sigma_{i}+\chi(\sigma_{i}>0)+\chi(i\in Z))_{i\in[1,|\sigma|]}$ the child of $\sigma$ corresponding to the subset of zeros $Z$ to which 1 is added. It should be apparent that $\sigma$ has $2^{|\mathbf{Zeros}(\sigma)|}$ children. For instance, the sequence $\sigma=(0,1,1,3,0,2,5)$ satisfies $\mathbf{Zeros}(\sigma)=\{1,5\}$, and its four children are
$\mathbf{child}(\sigma,\emptyset)$ $=(0,0,2,2,4,0,3,6),$
$\mathbf{child}(\sigma,\{1\})$ $=(0,1,2,2,4,0,3,6),$
$\mathbf{child}(\sigma,\{5\})$ $=(0,0,2,2,4,1,3,6),$
$\mathbf{child}(\sigma,\{1,5\})$ $=(0,1,2,2,4,1,3,6).$
Up to a change of labels, the generating tree for inversion sequences growing on the left is described by the succession rule $\Omega_{\text{Left}}$ below, where the new label $k$ corresponds to the number of zeros in a sequence. This generating tree is represented with both labellings on Figure 1.
$\Omega_{\text{Left}}=\left\{\begin{array}[]{rclll}(0)\\ (k)&\leadsto&(i+1)^{\binom{k}{i}}&\text{for}&i\in[0,k]\\ \end{array}\right.$
[…]
Figure 1: First four levels of the generating tree growing on the left for all inversion sequences.
The simplest construction of inversion sequences is the generating tree growing sequences on the right, used in e.g. [Kotsireas_Mansour_Yildirim_2024] or [Testart2024, Section 2]. In comparison, growing sequences on the left is not the most practical way to construct the set of inversion sequences (it is not even obvious that the tree described by $\Omega_{\text{Left}}$ has $n!$ nodes at each level $n$). However, this construction becomes simpler for some classes of pattern-avoiding inversion sequences. For instance, [Corteel_Martinez_Savage_Weselcouch_2016] proves that the generating trees growing on the left for the classes $\mathcal{I}(101)$ and $\mathcal{I}(110)$ are both described by the succession rule
$\Omega_{\text{pCat}}=\left\{\begin{array}[]{rclll}(0)\\ (k)&\leadsto&(k+1)\\ &&(i)^{i}&\text{for}&i\in[1,k].\\ \end{array}\right.$
The generating tree growing on the left for inversion sequences lacks a property that is typically desired when working with objects that avoid patterns: it does not automatically restrict to a generating tree for families of pattern-avoiding inversion sequences. Specifically, restricting the tree to keep only sequences which avoid some patterns may disconnect the tree. This occurs because an inversion sequence which contains a pattern $\rho$ may have children which avoid $\rho$. For instance, the sequence $(0,0,2,1)$ avoids the pattern 010, but its parent sequence is $(0,1,0)$. However, we show in Proposition 1 below that this problem only occurs with patterns which contain several zeros. Even then, it can be worked around in some cases. For instance, we slightly modify our construction in Section 4 to describe a generating tree for the class $\mathcal{I}(010,102)$.
Proposition 1.
Let $\rho$ be a pattern of length $k\geqslant 1$ which contains a single zero, let $\sigma$ be an inversion sequence which contains an occurrence $(\sigma_{q_{i}})_{i\in[1,k]}$ of $\rho$, and let $\tau$ be a child of $\sigma$ in the generating tree growing on the left. Then $\tau$ contains the pattern $\rho$.
Proof.
Since $\rho$ contains a single zero, $(\sigma_{q_{i}})_{i\in[1,k]}$ contains at most one zero. It follows that $(\sigma_{q_{i}})_{i\in[1,k]}$ is order-isomorphic to $(\tau_{q_{i}+1})_{i\in[1,k]}$ (every nonzero value is increased by 1, and the hypothetical zero can either increase by 1 or remain a zero), hence $\tau$ contains the pattern $\rho$. ∎
3 Inversion sequences avoiding 201 and 210
For all $n\geqslant 0$ and $\sigma\in\mathcal{I}_{n}(201,210)$, let $L(\sigma)=\{i\in[1,n]\,:\,\forall j\leqslant i,\;\sigma_{j}=0\}$ be the set of leading zeros of $\sigma$, and let $R(\sigma)=\mathbf{Zeros}(\sigma)\backslash L(\sigma)$. Let $\ell(\sigma)=|L(\sigma)|$ be the number of leading zeros of $\sigma$. If $\ell(\sigma)<n$ (i.e. if $\sigma$ is not constant), then $\sigma_{\ell(\sigma)+1}$ is the value of the first nonzero entry of $\sigma$. Notice that $L(\sigma)$ (resp. $R(\sigma)$) is the set of zeros of $\sigma$ to the left (resp. right) of position $\ell(\sigma)+1$.
Proposition 2.
Let $\sigma\in\mathcal{I}_{n}(201,210)$, $Z\subseteq\mathbf{Zeros}(\sigma)$, and let $\tau=\mathbf{child}(\sigma,Z)\in\mathcal{I}_{n+1}$. Then $\tau$ avoids 201 and 210 if and only if $Z\cap R(\sigma)\in\{\emptyset,R(\sigma)\}$.
Proof.
Since the patterns 201 and 210 both begin with the value 2, the first value of an occurrence of either pattern in $\tau$ must be at least 2. For all $i\in[1,\ell(\sigma)+1]$, we have $\tau_{i}\leqslant 1$, therefore any occurrence of 201 or 210 in $\tau$ must be in $(\tau_{i})_{i\in[\ell(\sigma)+2,n+1]}$. If $Z\cap R(\sigma)\in\{\emptyset,R(\sigma)\}$, then $(\tau_{i})_{i\in[\ell(\sigma)+2,n+1]}$ is order-isomorphic to $(\sigma_{i})_{i\in[\ell(\sigma)+1,n]}$, which avoids both patterns.
Assume $Z\cap R(\sigma)\notin\{\emptyset,R(\sigma)\}$. Let $i,j\in R(\sigma)$ be such that $i\in Z$ and $j\notin Z$. In particular, $\tau_{\ell(\sigma)+2}\geqslant 2$, $\tau_{i+1}=1$, and $\tau_{j+1}=0$. If $i<j$, then $(\tau_{\ell(\sigma)+2},\tau_{i+1},\tau_{j+1})$ is an occurrence of 210. If $j<i$, then $(\tau_{\ell(\sigma)+2},\tau_{j+1},\tau_{i+1})$ is an occurrence of 201. ∎
The proposition above establishes that in the generating tree growing on the left for $\mathcal{I}(201,210)$, the children of $\sigma$ are the sequences such that the subset $Z$ of zeros to which the value 1 is added must contain either all or none of $R(\sigma)$. Any subset of $L(\sigma)$ may intersect with $Z$ without creating an occurrence of 201 or 210. It follows that the number of children of $\sigma$ is $2^{\ell(\sigma)+\chi(R(\sigma)\neq\emptyset)}$.
For all $\sigma\in\mathcal{I}(201,210)$, let $s(\sigma)=\ell(\sigma)+\chi(R(\sigma)\neq\emptyset)$.
Lemma 3.
The generating tree growing on the left for $\mathcal{I}(201,210)$ is described by the succession rule
$\Omega_{\{201,210\}}=\left\{\begin{array}[]{rclll}(0)\\ (s)&\leadsto&(i)&\text{for}&i\in[1,s+1]\\ &&(i+1)^{2^{s-i}-1}&\text{for}&i\in[1,s-1].\end{array}\right.$
Proof.
We show that the succession rule $\Omega_{\{201,210\}}$ is obtained by replacing each sequence $\sigma$ by the label $s(\sigma)$ in the generating tree growing on the left for $\mathcal{I}(201,210)$. First, the sequence $\varepsilon$ corresponding to the root has label $s(\varepsilon)=0$.
Let $\sigma\in\mathcal{I}(201,210)$. We know from Proposition 2 that the set of children of $\sigma$ in the generating tree growing on the left for $\mathcal{I}(201,210)$ is
$\{\mathbf{child}(\sigma,Z)\;:\;Z\subseteq L(\sigma)\}\cup\{\mathbf{child}(\sigma,Z^{\prime}\sqcup R(\sigma))\;:\;Z^{\prime}\subseteq L(\sigma)\}.$
Note that the above union is disjoint if $R(\sigma)$ is nonempty, and both sets in the union are equal if $R(\sigma)$ is empty. Let $Z$ be such that $\tau=\mathbf{child}(\sigma,Z)\in\mathcal{I}(201,210)$ (i.e. $Z\subseteq L(\sigma)$, or $Z=Z^{\prime}\sqcup R(\sigma)$ for some $Z^{\prime}\subseteq L(\sigma)$). We distinguish several cases:
  • If $R(\sigma)=\emptyset$, then $s(\sigma)=\ell(\sigma)$.
    • If $Z=[i,\ell(\sigma)]$ for some $i\in[1,\ell(\sigma)+1]$, then $L(\tau)=[1,i]$ and $R(\tau)=\emptyset$, therefore $s(\tau)=i$.
    • Otherwise $R(\tau)\neq\emptyset$. We have $s(\tau)=i+1$, where $i=\ell(\tau)=\mathbf{min}(Z)\in[1,\ell(\sigma)-1]$. There are $12^{s(\sigma)-i}-$ 1 possible sets $Z$ in that case (that is all sets of the form $Z=\{i\}\sqcup Z^{\prime}$ for some $Z^{\prime}\subsetneq[i+1,\ell(\sigma)]$).
  • If $R(\sigma)\neq\emptyset$, then $s(\sigma)=\ell(\sigma)+1$.
    • If $Z=\emptyset$, then $L(\tau)=[1,\ell(\sigma)+1]$ and $R(\tau)=\{i+1\;|\;i\in R(\sigma)\}$, therefore $s(\tau)=s(\sigma)+1$.
    • If $Z=\mathbf{Zeros}(\sigma)\cap[i,n]$ for some $i\in[1,\ell(\sigma)+1]$, then $L(\tau)=[1,i]$ and $R(\tau)=\emptyset$, therefore $s(\tau)=i$.
    • Otherwise, $Z$ is not empty nor consists of a set of final zeros of $\sigma$. This implies that $R(\tau)\neq\emptyset$, and that $s(\tau)=i+1$, where $i=\ell(\tau)=\mathbf{min}(Z)\in[1,\ell(\sigma)]$. There are $12^{s(\sigma)-i}-$ 1 possible sets $Z$ in that case: either $Z=\{i\}\sqcup Z^{\prime}$ for some $Z^{\prime}\subseteq[i+1,\ell(\sigma)]$, or $Z=\{i\}\sqcup Z^{\prime}\sqcup R(\sigma)$ for some $Z^{\prime}\subsetneq[i+1,\ell(\sigma)]$). ∎
This succession rule gives a new, simple way of computing the generating function of $\mathcal{I}(201,210)$. For all $n,k\geqslant 0$, let $\mathfrak{a}_{n,k}=|\{\sigma\in\mathcal{I}_{n}(201,210)\;:\;k=s(\sigma)\}|$. Let $A(t,u)=\sum_{n,k\geqslant 0}\mathfrak{a}_{n,k}t^{n}u^{k}$.
Theorem 4.
The generating function of $\mathcal{I}(201,210)$ is $A(t,1)=\frac{2-t-t\sqrt{1-8t}}{4t^{2}-4t+2}$.
Proof.
The succession rule $\Omega_{\{201,210\}}$ gives the following equation:
$A(t,u)$ $=1+\sum_{n,k\geqslant 0}\mathfrak{a}_{n,k}t^{n+1}\left(\sum_{i=1}^{k+1}u^{i}+\sum_{i=1}^{k-1}(2^{k-i}-1)u^{i+1}\right)$
$=1+\sum_{n,k\geqslant 0}\mathfrak{a}_{n,k}t^{n+1}\left(u+\sum_{i=1}^{k}2^{k-i}u^{i+1}\right)$
$=1+\sum_{n,k\geqslant 0}\mathfrak{a}_{n,k}t^{n+1}\left(u+2^{k}u\sum_{i=1}^{k}\left(\frac{u}{2}\right)^{i}\right)$
$=1+\sum_{n,k\geqslant 0}\mathfrak{a}_{n,k}t^{n+1}\left(u+u^{2}\frac{2^{k}-u^{k}}{2-u}\right)$
$=1+tuA(t,1)+\frac{tu^{2}}{2-u}(A(t,2)-A(t,u)).$
This can be rewritten as
$\left(1+\frac{tu^{2}}{2-u}\right)A(t,u)=\frac{tu^{2}}{2-u}A(t,2)+tuA(t,1)+1.$ (1)
We apply the kernel method. The kernel in (1) is $1+\frac{tu^{2}}{2-u}$. To cancel the kernel, we substitute $u$ with $U(t)=\frac{1-\sqrt{1-8t}}{2t}$. This yields the equation
$A(t,2)=tU(t)A(t,1)+1.$ (2)
Evaluating (1) for $u=1$ gives
$A(t,1)=tA(t,2)+1.$
Replacing $A(t,2)$ by its expression from (2), we obtain
$A(t,1)=t^{2}U(t)A(t,1)+t+1.$
This gives an expression for $A(t,1)$:
$A(t,1)$ $=\frac{t+1}{1-t^{2}U(t)}=\frac{2-t-t\sqrt{1-8t}}{4t^{2}-4t+2}.\qed$
4 Inversion sequences avoiding 010 and 102
Proposition 5.
Let $\sigma\in\mathcal{I}(010)$. Then
  1. 1.
    all entries of value 0 in $\sigma$ must be consecutive and at the beginning of $\sigma$,
  2. 2.
    all entries of value 1 in $\sigma$ must be consecutive.
Proof.
  1. 1.
    Assume for the sake of contradiction that $\sigma$ contains two non-consecutive entries of value 0. In other words, there exists a subsequence $(\sigma_{i},\sigma_{j},\sigma_{k})$ such that $i<j<k$, $\sigma_{i}=\sigma_{k}=0$, and $\sigma_{j}\neq 0$. In particular, $\sigma_{j}\geqslant 1$, therefore $(\sigma_{i},\sigma_{j},\sigma_{k})$ is an occurrence of the pattern 010.
    We showed that all entries of value 0 in $\sigma$ must be consecutive. Since any nonempty inversion sequence $\sigma$ satisfies $\sigma_{1}=0$, the consecutive entries of value 0 must be at the beginning of $\sigma$.
  2. 2.
    Assume for the sake of contradiction that $\sigma$ contains two non-consecutive entries of value 1. In other words, there exists a subsequence $(\sigma_{i},\sigma_{j},\sigma_{k})$ such that $i<j<k$, $\sigma_{i}=\sigma_{k}=1$, and $\sigma_{j}\neq 1$. We showed that all entries of value 0 are consecutive and at the beginning of $\sigma$, hence $\sigma_{j}\neq 0$. It follows that $\sigma_{j}\geqslant 2$, therefore $(\sigma_{i},\sigma_{j},\sigma_{k})$ is an occurrence of the pattern 010. ∎
Recall that the generating tree growing on the left construction as we described it in Section 2 cannot be applied to the class $\mathcal{I}(010,102)$, due to the avoidance of the pattern 010. More precisely, applying this construction to every sequence in $\mathcal{I}_{n}(010,102)$ generates every sequence of $\mathcal{I}_{n+1}(010,102)$ in which the value 1 does not appear to the right of a greater value. In order to generate every sequence of $\mathcal{I}(010,102)$, we need to complete this construction with new steps which enable the value 1 to appear at any position (while avoiding the patterns 010 and 102).
The most obvious solution is to simply insert new entries of value 1. Usually, inserting a new entry before the end of an inversion sequence results in a construction which is not surjective. Indeed, if a sequence $\tau$ is obtained by inserting a new entry of value 1 at position $i$ in an inversion sequence $\sigma\in\mathcal{I}_{n}$, then for all $j\in[i+1,n+1]$, we have $\tau_{j}=\sigma_{j-1}<j-1$. In particular, a sequence $\tau$ which contains a saturated entry $\tau_{j}=j-1$ for some $j>i$ cannot be generated this way. However in our case, we are interested in generating a sequence $\tau$ which avoids 102, and in which the value 1 at position $i$ appears to the right of a value $k>1$. In particular, for all $j>i$, $\tau$ must satisfy $\tau_{j}\leqslant k<i-1<j-1$, so the surjectivity problem disappears.
It can be shown that the construction described above indeed defines a generating tree for $\mathcal{I}(010,102)$. In this tree, the children of a sequence are obtained by either applying the same steps as for the generating tree growing on the left, or inserting a new entry of value 1 (to the right of a greater value). In other words, the parent of a nonempty sequence is obtained by
  • removing an entry of value 1, if such an entry appears to the right of a greater value (it does not matter which entry of value 1 is removed since they are all consecutive by Proposition 5);
  • removing the first zero and subtracting 1 from all nonzero entries, otherwise.
We introduce a second small change in the construction, to make it more convenient later. If a nonempty sequence contains an entry of value 1 with no greater value to its left (i.e. the sequence begins with a prefix of zeros, followed by a 1), then its parent is now obtained by subtracting 1 from the leftmost entry of value 1 only [^3] When exploring the ascendants of a sequence of this form, its parent in our original definition will eventually appear, after all entries of value 1 have been turned into zeros. We have simply placed some intermediate sequences on that path.. Otherwise, its parent is determined as described just above. The corresponding tree is represented in Figure 2. It should be noted that this tree does not quite conform to the usual definition of a generating tree, as some nodes are on the same level as their parent.
The article [JumpingGT] introduces the notion of ”jumping” generating trees, in which each edge has a (bounded) positive integer length. If a node at level $n$ has a child through an edge of length $k$, then this child is at level $n+k$ (in a classical generating tree, all edges have length $1$). The tree in Figure 2 could be seen as a jumping generating tree having edges of length 0 or 1.
This generating tree allows us to compute the generating function of $\mathcal{I}(010,102)$. However, this result is easier to obtain by restricting our study to the subset of sequences of $\mathcal{I}(010,102)$ which do not contain the value 1. Due to this restriction, the construction presented in the remainder of this section jumps arbitrarily large lengths, so we shall no longer speak of it as a generating tree.
$\varepsilon$0000000000000000001000020011011100022001000200003000210012000230112011002200033002210021201220100200030000400031001300024011300210003200211012002300034002310123
Figure 2: Part of a generating tree for $\mathcal{I}(010,102)$. Sequences containing the value 1 are colored in gray.
[…]
For all $\sigma\in\mathcal{I}$, let $\mathbf{push}(\sigma)=0\cdot(\sigma_{i}+\chi(\sigma_{i}>0))_{i\in[0,n]}$ be the sequence obtained by inserting a zero at the beginning of $\sigma$, and adding 1 to the value of all nonzero entries.
Proposition 6.
The image $\mathbf{push}(\mathcal{I}(010,102))$ is the set $\mathfrak{A}\backslash\{\varepsilon\}$.
Proof.
Let $\sigma\in\mathcal{I}(010,102)$, and let $\tau=\mathbf{push}(\sigma)$. By construction, $\tau$ is an inversion sequence, $\tau$ does not contain the value $1$, and $|\tau|>|\sigma|$ so $\tau$ cannot be empty. Let us now show that $\tau$ avoids 010 and 102. The sequence $(\tau_{i})_{i\in[2,n+1]}$ is order-isomorphic to $\sigma$, which avoids 010 and 102. This means that any occurrence of either pattern in $\tau$ must involve the first entry of $\tau$. Since $\tau_{1}=0$, it cannot be part of an occurrence of 102. If $\tau$ contained an occurrence $(\tau_{1},\tau_{i},\tau_{j})$ of 010 for some $i<j$, then $\sigma$ would contain a zero $\sigma_{j-1}$ after a nonzero value $\sigma_{i-1}$, which is impossible by Proposition 5.
Let $\tau\in\mathfrak{A}\backslash\{\epsilon\}$, and let $\sigma$ be the sequence obtained by removing the first zero of $\tau$ and subtracting 1 from the value of all nonzero entries. Then $\sigma\in\mathcal{I}(010,102)$, and $\mathbf{push}(\sigma)=\tau$. ∎
Corollary 7.
For all $n\geqslant 0$, $|\mathcal{I}_{n}(010,102)|=|\mathfrak{A}_{n+1}|$.
Proof.
Let $n\geqslant 0$. The function $\mathbf{push}$ is injective, so it describes a bijection between $\mathcal{I}_{n}(010,102)$ and $\mathbf{push}(\mathcal{I}_{n}(010,102))$. By construction, the image under $\mathbf{push}$ of a sequence of length $n$ is a sequence of length $n+1$, so it follows from Proposition 6 that $\mathbf{push}(\mathcal{I}_{n}(010,102))=\mathfrak{A}_{n+1}$. ∎
For all $n\geqslant 0$ and $\sigma\in\mathcal{I}_{n}(010,102)$, let
$\mathbf{Sites}(\sigma)=\{i\in[2,n+1]\;:\;(\sigma_{j})_{j\in[1,i-1]}\cdot 1\cdot(\sigma_{j})_{j\in[i,n]}\in\mathcal{I}_{n+1}(010,102)\}$
be the set of positions, called active sites, where inserting the value 1 in $\sigma$ does not create an occurrence of 010 or 102. Note that we do not allow the value 1 to be inserted at the beginning of $\sigma$ (i.e. at position 1), since the result would not be an inversion sequence.
If $\sigma$ already contains the value 1, then a new entry of value 1 can only be inserted next to the current entries of value 1, by Proposition 5, and we have little interest in the active sites of $\sigma$ in that case. We next characterize the active sites of $\sigma$ when $\sigma$ does not contain the value 1.
Proposition 8.
Let $n\geqslant 1$, $\sigma\in\mathfrak{A}_{n}$, and $z=|\mathbf{Zeros}(\sigma)|$. Then
$\mathbf{Sites}(\sigma)=\{i\in[z+1,n+1]\;:\;\mathbf{min}((\sigma_{j})_{j\in[z+1,i-1]})\geqslant\mathbf{max}((\sigma_{j})_{j\in[i,n]})\}.$
Proof.
The value 1 cannot be inserted at positions $[2,z]$, since it would create a subsequence $(0,1,0)$. Inserting the value 1 at any other position (i.e. in $[z+1,n+1])$ cannot create an occurrence of 010, since there are no zeros to the right and no other entry of value $1$ in $\sigma$. It follows that $\mathbf{Sites}(\sigma)$ is the subset of $[z+1,n+1]$ of positions where inserting the value 1 does not create an occurrence of 102. Note that all values of $(\sigma_{i})_{i\in[z+1,n]}$ are greater than 1, so the new entry of value 1 could only take the role of the ’0’ of the pattern 102. In other words, $i\in[z+1,n+1]$ is an active site of $\sigma$ if and only if no value of $(\sigma_{j})_{j\in[z+1,i-1]}$ is less than a value of $(\sigma_{j})_{j\in[i,n]}$. ∎
Let $\mathfrak{B}=\mathfrak{A}\backslash\{(0)^{n}\;:\;n\geqslant 0\}$ be the subset of non-constant sequences in $\mathfrak{A}$. Equivalently, $\mathfrak{B}$ is the set of inversion sequences which avoid 010 and 102, do not contain the value 1, and contain a value greater than 1. For all $n,z,s\geqslant 0$, let $\mathfrak{B}_{n}=\{\sigma\in\mathfrak{B}\;:\;|\sigma|=n\}$, let $\mathfrak{B}_{n,z}=\{\sigma\in\mathfrak{B}_{n}\;:\;|\mathbf{Zeros}(\sigma)|=z\}$, and let $\mathfrak{B}_{n,z,s}=\{\sigma\in\mathfrak{B}_{n,z}\;:\;|\mathbf{Sites}(\sigma)|=s\}$. Let also $\mathfrak{b}_{n}=|\mathfrak{B}_{n}|$, $\mathfrak{b}_{n,z}=|\mathfrak{B}_{n,z}|$, and $\mathfrak{b}_{n,z,s}=|\mathfrak{B}_{n,z,s}|$.
Lemma 9.
Let $n,z,s\geqslant 0$. If $n\geqslant 3$, $z\in[2,n-1]$, and $s\in[2,n-z+1]$, then
$\mathfrak{b}_{n,z,s}$ $=\mathfrak{b}_{n-1,z-1,s}+\chi(z+s=n+1)+\chi(s=2)\cdot\left(\sum_{z^{\prime}\geqslant z}\mathfrak{b}_{n-1,z^{\prime}}\right)$
$+\left(\sum_{k\geqslant 1}\mathfrak{b}_{n-k-1,z-1,s-k}\right)+\chi(s\geqslant 3)\cdot\left(\sum_{n^{\prime}\leqslant n-2}\sum_{s^{\prime}\geqslant s}\mathfrak{b}_{n^{\prime},z-1,s^{\prime}}\right).$
Otherwise, $\mathfrak{b}_{n,z,s}=0$.
Proof.
Let $\sigma\in\mathfrak{B}_{n,z,s}$. By definition, $\sigma$ is not constant, so we have $z<n$. By definition, $\sigma$ cannot contain the value 1, hence $\sigma$ begins by $(0,0)$, and $z\geqslant 2$. This implies that $n\geqslant 3$. By Proposition 8, $\mathbf{Sites}(\sigma)\subseteq[z+1,n+1]$, and $z+1$ and $n+1$ are always active sites of $\sigma$, so $s\in[2,n-z+1]$.
Let $\sigma^{\prime}=\mathbf{push}^{-1}(\sigma)\in\mathcal{I}_{n-1}(010,102)$ be the sequence obtained by removing the first zero from $\sigma$, and subtracting 1 from every nonzero entry.
  • If $\sigma^{\prime}$ does not contain the value 1, then $\sigma^{\prime}\in\mathfrak{b}_{n-1,z-1,s}$.
  • Otherwise, $\sigma^{\prime}$ has entries of value 1, and they are all consecutive by Proposition 5.
    • If $\sigma^{\prime}_{z}=1$ (i.e. the factor of ones appears just after the factor of zeros),
        If $\sigma^{\prime}_{n-1}=1$, then $\sigma^{\prime}=(0)^{z-1}\cdot(1)^{n-z}$ by Proposition 5, so $\sigma=(0)^{z}\cdot(2)^{n-z}$. In that case, $\mathbf{Sites}(\sigma)=[z+1,n+1]$, so $z+s=n+1$.
        Otherwise, $\sigma^{\prime}_{n-1}>1$. In that case, we have $\sigma_{z+1}=2$ and $\sigma_{n}>2$, hence inserting the value 1 at any position in $[z+2,n]$ would create an occurrence of 102. This means that $\mathbf{Sites}(\sigma)=\{z+1,n+1\}$, so $s=2$.
        Since the factor of ones of $\sigma^{\prime}$ is consecutive to its factor of zeros, replacing all the ones by zeros yields a sequence $\tau$ which still avoids 010 and 102. Notice that $\tau$ does not contain the value 1, but contains a value greater than 1, so $\tau\in\mathfrak{B}_{n-1}$. By construction, the number $z^{\prime}$ of zeros of $\tau$ is at least $z$.
    • If $\sigma^{\prime}_{z}>1$ and $\sigma^{\prime}_{n-1}=1$, let $k$ be the number of entries of value $1$ in $\sigma^{\prime}$, and let $\tau=(\sigma^{\prime}_{i})_{i\in[1,n-k-1]}$ be the sequence obtained by removing all entries of value 1 from $\sigma^{\prime}$. In that case, the rightmost $k$ positions $[n-k+2,n+1]$ are active sites of $\sigma$ which are removed in $\tau$, so $\tau\in\mathfrak{B}_{n-k-1,z-1,s-k}$.
    • If $\sigma^{\prime}_{z}>1$ and $\sigma^{\prime}_{n-1}>1$, let $p\in[z+2,n-1]$ be the position of the leftmost entry of value 2 in $\sigma$ (such an entry exists because $\sigma^{\prime}$ contains the value 1). Then $\{z+1,p,n+1\}\subseteq\mathbf{Sites}(\sigma)$, so $s\geqslant 3$.
      Let $\tau$ be the sequence obtained by removing all entries of value 1 from $\sigma^{\prime}$. Let $n^{\prime}=|\tau|$, and $s^{\prime}=|\mathbf{Sites}(\tau)|$, so that $\tau\in\mathfrak{B}_{n^{\prime},z-1,s^{\prime}}$. By construction, $n^{\prime}\leqslant n-2$. It can be seen that $\{i-1\;:\;i\in\mathbf{Sites}(\sigma)\backslash\{n+1\}\}\sqcup\{n^{\prime}+1\}\subseteq\mathbf{Sites}(\tau)$, so $s^{\prime}\geqslant s$.
To complete the proof, it remains to show that each of the five cases above induces a bijection between the set of sequences $\sigma$ and the set of sequences $\sigma^{\prime}$ or $\tau$. We describe the inverse bijection only for the final case, since the previous cases are simpler.
Let $n\geqslant 3$, $z\geqslant 2$, $s\geqslant 3$, $n^{\prime}\leqslant n-2$, $s^{\prime}\geqslant s$, and let $\tau\in\mathfrak{B}_{n^{\prime},z-1,s^{\prime}}$. Let also $p_{1}<p_{2}<\dots<p_{s^{\prime}}$ be the active sites of $\tau$. Let $\sigma^{\prime}$ be the sequence obtained by inserting $n-n^{\prime}-1$ entries of value 1 at position $p_{s-1}$, and let $\sigma=\mathbf{push}(\sigma^{\prime})$. By construction, $|\sigma|=n$ and $|\mathbf{Zeros}(\sigma)|=z$. Notice that inserting the value 1 in $\sigma$ at any position in $[p_{s-1}+2,n]$ would create an occurrence of 102, since $\sigma_{p_{s-1}+1}=2$ and $\sigma_{n}>2$. It follows that the active sites of $\sigma$ are $\{p_{i}+1\;:\;i\in[1,s-1]\}\sqcup\{n+1\}$, so $|\mathbf{Sites}(\sigma)|=s$. ∎
Let $F(t)=\sum_{n\geqslant 0}|\mathcal{I}_{n}(010,102)|t^{n}$ be the ordinary generating function of $\mathcal{I}(010,102)$. Let also $B(t,u,v)=\sum_{n,z,s\geqslant 0}\mathfrak{b}_{n,z,s}t^{n}u^{z}v^{s}$.
Theorem 10.
The function $F(t)$ is algebraic with minimal polynomial
$(t^{5}-3t^{4}+4t^{3}-3t^{2}+t)F(t)^{3}+(4t^{4}-8t^{3}+6t^{2}-2t)F(t)^{2}$
$+(-t^{4}+8t^{3}-11t^{2}+6t-1)F(t)-2t^{3}+5t^{2}-4t+1.$
Proof.
We know from the definition of $\mathfrak{B}$ that $|\mathfrak{A}_{n}|=\mathfrak{b}_{n}+1$ for all $n\geqslant 0$, so it follows from Corollary 7 that
$F(t)=\frac{1}{t}\left(B(t,1,1)+\frac{1}{1-t}-1\right),$
and we proceed by computing $B(t,1,1)$.
The formula of Lemma 9 is equivalent to the functional equation
$B(t,u,v)$ $=tuB(t,u,v)+\frac{t^{3}u^{2}v^{2}}{(1-tu)(1-tv)}+tv^{2}\left(\frac{u^{2}B(t,1,1)-uB(t,u,1)}{1-u}\right)$
$+t^{2}uv\left(\frac{B(t,u,v)}{1-tv}\right)+t^{2}u\left(\frac{v^{3}B(t,u,1)-vB(t,u,v)}{(1-t)(1-v)}\right).$
This can be rewritten as
$K_{1}(t,u,v)B(t,u,v)=K_{2}(t,u,v)B(t,u,1)+\frac{tu^{2}v^{2}}{1-u}B(t,1,1)+\frac{t^{3}u^{2}v^{2}}{(1-tu)(1-tv)},$ (3)
where $K_{1}$ and $K_{2}$ are rational functions of $(t,u,v)$:
$K_{1}(t,u,v)$ $=\frac{-t^{3}uv^{2}-t^{2}v^{2}+t^{2}u+t^{2}v+tuv+tv^{2}-tu-t-v+1}{(1-t)(1-v)(1-tv)},$
$K_{2}(t,u,v)$ $=\frac{tuv^{2}(-tuv+t+v-1)}{(1-t)(1-u)(1-v)}.$
We proceed with the kernel method. The kernel in (3) is $K_{1}(t,u,v)$. This kernel can be cancelled by substituting $v$ for a formal power series $V(t,u)$ which is algebraic with minimal polynomial
$(t^{3}u+t^{2}-t)V(t,u)^{2}+(-t^{2}-tu+1)V(t,u)-t^{2}u+tu+t-1.$
An explicit expression for $V(t,u)$ is
$\frac{t^{2}+tu-1+\sqrt{4t^{5}u^{2}-4t^{4}u^{2}+t^{4}-2t^{3}u+t^{2}u^{2}-4t^{3}+4t^{2}u+6t^{2}-2tu-4t+1}}{2t(t^{2}u+t-1)}.$
Substituting $v$ for $V(t,u)$ in (3) yields the equation
$K_{2}(t,u,V(t,u))B(t,u,1)=-\frac{tu^{2}V(t,u)^{2}}{1-u}B(t,1,1)-\frac{t^{3}u^{2}V(t,u)^{2}}{(1-tu)(1-tV(t,u))}.$ (4)
The kernel in (4) is $K_{2}(t,u,V(t,u))$, and it can be cancelled by substituting $u$ for a formal power series $U(t)$ which is algebraic with minimal polynomial
$t^{2}U(t)^{3}-2tU(t)^{2}+(-t^{3}+2t^{2}+1)U(t)-t^{2}+t-1.$
Substituting $u$ for $U(t)$ in (4), we obtain
$B(t,1,1)=-\frac{(1-U(t))t^{3}U(t)^{2}V(t,U(t))^{2}}{tU(t)^{2}V(t,U(t))^{2}(1-tU(t))(1-tV(t,U(t)))}.$
From this equation, and the minimal polynomials of $U$ and $V$, we can compute [^4] We use the eliminate command in Maple. the minimal polynomial of $B(t,1,1)$:
$(t^{3}-2t^{2}+2t-1)B(t,1,1)^{3}+(t^{3}-t^{2}-t)B(t,1,1)^{2}+(-t^{4}+2t^{3}-4t^{2}+t)B(t,1,1)-t^{4}.$
[…]
$(t^{5}-3t^{4}+4t^{3}-3t^{2}+t)F(t)^{3}+(4t^{4}-8t^{3}+6t^{2}-2t)F(t)^{2}$
$+(-t^{4}+8t^{3}-11t^{2}+6t-1)F(t)-2t^{3}+5t^{2}-4t+1.\qed$
Acknowledgements
We thank Mathilde Bouvel and Emmanuel Jeandel for their comments which improved the presentation of this work.
[…]